what is the y intercept of the line that passes through point (-2,1) and is perpendicular to line y=-3x+1
Answer by TimothyLamb(4379)

perp lines have negative reciprocal slopes
---
a line perp to:
y = -3x + 1
has slope: 1/3
---
use the point-slope form: y - y1 = m(x - x1)
with point: (-2,1)
y - 1 = (1/3)(x + 2)
---
y = (1/3)x + 2/3 + 1
y = (1/3)x + 2/3 + 3/3
y = (1/3)x + 5/3
Answer:
b = 5/3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php